# A posteriori error estimation for an interior penalty virtual element method of Kirchhoff plates

### v2

@@ Abstract
A residual-type a posteriori error estimation is developed for an interior penalty virtual element method (IPVEM) to solve a Kirchhoff plate bending problem with inhomogeneous boundary value conditions. The computable error estimator is incorporated. We derive the reliability and efficiency of the a posteriori error bound by constructing an enriching operator and establishing some related error estimates that align with interior penalty finite element methods. As an outcome of the error estimator, an adaptive VEM is introduced by means of the mesh refinement strategy with the one-hanging-node rule. Numerical results on several benchmark tests confirm the robustness of the proposed error estimator and show the efficiency of the resulting adaptive VEM.
@@ 1 Introduction
Virtual element methods (VEMs) generalize standard finite element methods to accommodate general polytopal meshes. Initially proposed and analyzed in [3], with subsequent foundational work in [1, 4], VEMs offer advantages in handling complex geometries and high-regularity solutions compared to traditional finite element methods. For the plate bending problem or the biharmonic equation, Brezzi and Marini introduced an $H^{2}$-conforming VEM in [11], establishing convergence results in the $H^{2}$ norm. Chinosi and Marini further improved the method in [17] by deriving optimal order error estimates in the $L^{2}$ and $H^{1}$ norms through VEM enhancement techniques. Nonconforming VEMs have also been developed, including the $C^{0}$-continuous $H^{2}$-nonconforming VEM presented in [28] and the Morley-type VEM discussed in [30]. Chen and Huang extended the approach to fully $H^{m}$-nonconforming VEMs for $2m$-th order problems in any dimension in [15]. Recently, the interior penalty virtual element method (IPVEM) has been explored in [29] for the biharmonic equation. Compared with the conforming or nonconforming VEMs for the biharmonic problem, the IPVEM has advantages in reducing the number of degrees of freedom (DoFs), which can be regarded as a combination of the virtual element space and discontinuous Galerkin scheme, since the resulting global discrete space is not $C^{0}$-continuous and an interior penalty formulation is adopted to enforce the continuity of the solution. This method was adapted for the fourth-order singular perturbation problem in [27], with some adaptations to the original space. These modifications encompass alterations in the definition of the $H^{1}$-type projection and the selection of the DoFs. In contrast, Ref. [18] employs the original IPVEM formulation from [29] to address the fourth-order singular perturbation problem (2.1), incorporating techniques from the modified Morley finite element method as described in [25].
@@
On the other hand, due to the large flexibility of the meshes in VEMs, the a posteriori error analysis of VEMs is well worth studying. For the fourth-order problem, a residual-based a posteriori error estimate for the $C^{1}$-continuous VEM was developed in [16], addressing both the reliability and efficiency of the error bound. Recently, the lowest-order nonconforming or Morley-type virtual element has been applied to the biharmonic equation, with a residual-based a posteriori error estimator derived in [13]. It’s worth noting that the a posteriori error analysis in [13] circumvents any trace of second derivatives by using some computable conforming companion operator, which significantly reduces the computational complexity of the error estimators. This reduction is facilitated by the special property of the interpolation operator for the Morley virtual element (see Lemma 5.1).
In this paper, we consider the a posteriori error estimation for an interior penalty virtual element method of Kirchhoff plates. We provide a detailed description for implementing the IPVEM with inhomogeneous boundary conditions. For the sake of the construction of error estimators, unlike the previous interior penalty formulations in [29, 18], we substitute the $H^{1}$-elliptic projection operator in the jump and average terms for second-order derivatives with an $H^{2}$-elliptic projection operator. Following the idea in [21], we construct an enriching operator to establish the reliability and efficiency of the a posteriori error bound. This enriching operator inherits the beneficial properties of the operators described in [13] for the lowest-order Morley virtual element and is applicable to arbitrary-order cases.
@@
The remainder of the paper is structured as follows. We begin by introducing the continuous variational problem and presenting some useful results in VEM analysis in Section 2. Section 3 details the interior penalty virtual element method, including the virtual element space, degrees of freedom, and the discrete bilinear form, with a particular focus on addressing inhomogeneous boundary value conditions not covered in previous references. In Section 4, we construct an enriching operator and present the corresponding error estimates. Section 5 develops the a posteriori analysis for the IPVEM, providing both upper and lower bounds. Meanwhile, we make a comparison with the analysis of Morley-type virtual elements in [13]. An adaptive IPVEM is presented in Section 6 with several benchmark tests performed to illustrate the robustness of the residual-type a posteriori error estimator and demonstrate the efficiency of the adaptive VEM.
@@ 2 The continuous variational problem
$\left\{\begin{array}[]{ll}\Delta^{2}u=f&{\rm in}~{}~{}\Omega,\\ u=g_{D},\quad\dfrac{\partial u}{\partial\bm{n}}=g_{N}&{\rm on}~{}~{}\partial\Omega.\end{array}\right.$ (2.1)
where $\bm{n}=(n_{1},n_{2})$ is the unit outer normal to $\partial\Omega$, $\Delta$ is the standard Laplacian operator, and $g_{D}\in H^{3/2}(\partial\Omega)$ and $g_{N}\in H^{1/2}(\partial\Omega)$ are the Dirichlet data. Throughout the article we assume $u\in H^{k+1}(\Omega)$ with $k\geq 2$
@@
$\sum_{e\in\mathcal{E}_{h}}\int_{e}[vw]\mathrm{d}s=\sum_{e\in\mathcal{E}_{h}^{0}}\int_{e}[v]\{w\}\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}}\int_{e}\{v\}[w]\mathrm{d}s,$ (2.2)
which will be used frequently later. Moreover, for any two quantities $a$ and $b$, “$a\lesssim b$” indicates “$a\leq Cb$” with the constant $C$ independent of the mesh size $h_{K}$, and “$a\eqsim b$” abbreviates “$a\lesssim b\lesssim a$”.
@@
where $a(u,v)=(\nabla^{2}u,\nabla^{2}v)$.
To keep the flexibility of meshes, we will work on meshes satisfying the following assumption (cf. [10, 14]):
@@
As shown in [14], this condition covers the usual geometric assumptions frequently used in the context of VEMs. According to the standard Dupont-Scott theory (cf. [7]), for all $v\in H^{l}(K)$ ($0\leq l\leq k$) there exists a certain $q\in\mathbb{P}_{l-1}(K)$ such that
@@
The following trace inequality [7] and the Sobolev inequality [6] are very useful for our forthcoming analysis.
Lemma 2.1 (Trace inequality).
@@
$\|v\|_{0,\partial K}\lesssim h_{K}^{1/2}|v|_{1,K}+h_{K}^{-1/2}\|v\|_{0,K},\quad v\in H^{1}(K).$ (2.5)
Lemma 2.2 (Sobolev inequality).
For any $K\in\mathcal{T}_{h}$, it holds
$\|v\|_{\infty,K}\lesssim h_{K}|v|_{2,K}+|v|_{1,K}+h_{K}^{-1}\|v\|_{0,K},\quad v\in H^{2}(K).$ (2.6)
@@ 3 The interior penalty virtual element method
This section reviews the interior penalty virtual element space developed in [29] and details the treatment of inhomogeneous boundary value conditions for the IPVEM.
@@ 3.1   The IP virtual element spaces
$\widetilde{V}_{k+2}(K)=\left\{v\in H^{2}(K):\Delta^{2}v\in\mathbb{P}_{k}(K),v|_{e}\in\mathbb{P}_{k+2}(e),\partial_{\bm{n}}v|_{e}\in\mathbb{P}_{k+1}(e),~{}~{}e\subset\partial K\right\},\quad k\geq 2.$
@@
where $\mathcal{V}_{K}$ is the set of the vertices of $K$. By checking the right-hand side of the integration by parts formula:
$(\nabla v_{h},\nabla q)_{K}=-(v_{h},\Delta q)_{K}+\sum\limits_{e\subset\partial K}\int_{e}v_{h}\partial_{\bm{n}_{e}}q\,\mathrm{d}s.$
However, the goal of the IPVEM is to make $\Pi_{K}^{\nabla}v_{h}$ computable by only using the DoFs of $H^{1}$-conforming virtual element spaces given by (cf. [3, 1])
@@
To do so, Ref. [29] considered the approximation of the right-hand side by some numerical formula. In view of the $\mathbb{P}_{k}$ accuracy, namely $\Pi_{K}^{\nabla}v_{h}=v_{h}$ for $v_{h}\in\mathbb{P}_{k}(K)$, the modified $H^{1}$-projection is defined as
@@
With the help of $\Pi_{K}^{\nabla}$, we can employ the standard enhancement technique by substituting the redundant moments of $v$ with those of $\Pi_{K}^{\nabla}v$ [1]. The local interior penalty space is defined as
$V_{k}(K)=\{v\in W_{k}(K):\tilde{\chi}^{g}(v)=\tilde{\chi}^{g}(\Pi_{K}^{\nabla}v),\tilde{\chi}^{n}(v)=\tilde{\chi}^{n}(\Pi_{K}^{\nabla}v)\},$
where
$W_{k}(K):=\{v\in H^{2}(K):\Delta^{2}v\in\mathbb{P}_{k-2}(K),v|_{e}\in\mathbb{P}_{k+2}(e),\partial_{\bm{n}}v|_{e}\in\mathbb{P}_{k+1}(e),~{}~{}e\subset\partial K,$
$\hskip 113.81102pt(v,q)_{K}=(\Pi_{K}^{\nabla}v,q)_{K},\quad q\in\mathbb{M}_{k}(K)\backslash\mathbb{M}_{k-2}(K)\}.$
@@
Choosing $q=\Pi_{K}^{\Delta}v$ in (3.7), we can easily observe
$|\Pi_{K}^{\Delta}v|_{2,K}\lesssim|v|_{2,K},\qquad v\in H^{2}(K).$ (3.8)
3.2   The IPVEM with inhomogeneous boundary value conditions
Let $v\in H^{2}(K)$. The integration by parts gives
$(f,v)_{K}=(\Delta^{2}u,v)_{K}=(\nabla^{2}u,\nabla^{2}v)_{K}+\int_{\partial K}\Big{(}\frac{\partial\Delta u}{\partial\bm{n}_{K}}v+\frac{\partial^{2}u}{\partial\bm{n}_{K}\partial\bm{t}_{K}}\frac{\partial v}{\partial\bm{t}_{K}}\Big{)}\mathrm{d}s-\int_{\partial K}\frac{\partial^{2}u}{\partial\bm{n}_{K}^{2}}\frac{\partial v}{\partial\bm{n}_{K}}\mathrm{d}s.$
Using the formula in (2.2) and noting that $u\in H^{3}(\Omega)$, we get
yielding
To ensure both the symmetry and $C^{1}$-continuity of the discrete variational problem, we rewrite it as
$(f,v_{h})-\sum_{e\in\mathcal{E}_{h}^{\partial}}\int_{e}\Big{\{}\frac{\partial^{2}v_{h}}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}^{\partial}}\frac{\lambda_{e}^{3}}{|e|^{3}}\int_{e}[u][v_{h}]\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}^{\partial}}\frac{\lambda_{e}}{|e|}\int_{e}\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s$
where we have used the fact that $[u]_{e}=0$ and $[\frac{\partial u}{\partial\bm{n}_{e}}]_{e}=0$ for $e\in\mathcal{E}_{h}^{0}$.
Following the standard procedure in the VEMs, we define the discrete bilinear form as
$a_{h}^{K}(v,w)=a^{K}(\Pi_{K}^{\Delta}v,\Pi_{K}^{\Delta}w)+S^{K}(v-\Pi_{K}^{\Delta}v,w-\Pi_{K}^{\Delta}w),\quad v,w\in V_{k}(K),$

### v4

@@ Abstract
In this paper, we develop a residual-type a posteriori error estimation for an interior penalty virtual element method (IPVEM) for the Kirchhoff plate bending problem. Building on the work in [18], we adopt a modified discrete variational formulation that incorporates the $H^{1}$-elliptic projector in the jump and average terms. This allows us to simplify the numerical implementation by including the $H^{1}$-elliptic projector in the computable error estimators. We derive the reliability and efficiency of the a posteriori error bound by constructing an enriching operator and establishing some related error estimates that align with $C^{0}$-continuous interior penalty finite element methods. As observed in the a priori analysis, the interior penalty virtual elements exhibit similar behaviors to $C^{0}$-continuous elements despite its discontinuity. This observation extends to the a posteriori estimate since we do not need to account for the jumps of the function itself in the discrete scheme and the error estimators. As an outcome of the error estimator, an adaptive VEM is introduced by means of the mesh refinement strategy with the one-hanging-node rule. Numerical results from several benchmark tests confirm the robustness of the proposed error estimators and show the efficiency of the resulting adaptive VEM.
@@ 1 Introduction
Virtual element methods (VEMs) generalize standard finite element methods to accommodate general polytopal meshes. Initially proposed and analyzed in [3], with subsequent foundational work in [1, 4], VEMs offer advantages in handling complex geometries and high-regularity solutions compared to traditional finite element methods. For the plate bending problem or the biharmonic equation, Brezzi and Marini introduced an $H^{2}$-conforming VEM in [11], establishing convergence results in the $H^{2}$ norm. Chinosi and Marini further improved the method in [17] by deriving optimal order error estimates in the $L^{2}$ and $H^{1}$ norms through VEM enhancement techniques. Nonconforming VEMs have also been developed, including the $C^{0}$-continuous $H^{2}$-nonconforming VEM presented in [28] and the Morley-type VEM discussed in [31]. Chen and Huang extended the approach to fully $H^{m}$-nonconforming VEMs for $2m$-th order problems in any dimension in [15]. Recently, the interior penalty virtual element method (IPVEM) has been explored in [30] for the biharmonic equation. Compared with the conforming or nonconforming VEMs for the biharmonic problem, the IPVEM has advantages in reducing the number of degrees of freedom (DoFs), which can be regarded as a combination of the virtual element space and discontinuous Galerkin scheme. This method was adapted for the fourth-order singular perturbation problem in [27], with some adaptations to the original space. These modifications encompass alterations in the definition of the $H^{1}$-type projection and the selection of the DoFs. In contrast, Ref. [18] employs the original IPVEM formulation from [30] to address the fourth-order singular perturbation problem, incorporating techniques from the modified Morley finite element method as described in [25].
@@
Another point to be mentioned is that due to the large flexibility of the meshes in VEMs, there is no exceptional treatment for hanging nodes, which has great superiority in mesh refinement. Hence, a posteriori error analysis of VEMs is well worth studying. For the fourth-order problem, a residual-based a posteriori error estimate for the $C^{1}$-continuous VEM was developed in [16], addressing both the reliability and efficiency of the error bound. Recently, the lowest-order nonconforming or Morley-type virtual element has been applied to the biharmonic equation, with a residual-based a posteriori error estimator derived in [13]. It’s worth noting that the a posteriori error analysis in [13] circumvents any trace of second derivatives by using some computable conforming companion operator, which significantly reduces the computational complexity of the error estimators. This reduction is facilitated by the special property of the interpolation operator for the Morley virtual element (see Lemma 5.1). Recently, the medius error estimates for the interior penalty virtual element method for the biharmonic equation are proposed in [29].
In this paper, we consider a posteriori error estimation for the interior penalty virtual element method applied to Kirchhoff plate bending problems. Unlike the jump and average terms in [30], we include the $H^{1}$-elliptic projector $\Pi_{h}^{\nabla}$ in the penalty terms, as proposed in [18]. This modification enables the design of computable error estimators that incorporate the $H^{1}$-elliptic projector, thereby simplifying the implementation process. Following the approach in [21], we construct an enriching operator to establish the reliability of the a posteriori error estimators. We then develop a residual-type a posteriori error estimation for the modified IPVEM applied to Kirchhoff plates. It is worth noting that although the interior penalty virtual element method is discontinuous, as observed in the a priori error estimate, the interior penalty virtual elements exhibit behavior similar to $C^{0}$-continuous elements. This “quasi-$C^{0}$ continuity” can be extended to the a posteriori estimate since we do not need to account for the jumps of the function itself in the discrete scheme or the error estimators.
@@
The remainder of the paper is structured as follows. We begin by introducing the continuous variational problem and presenting some useful results in VEM analysis in Section 2. Section 3 describes the interior penalty virtual element method. We also incorporate inhomogeneous boundary conditions not covered in previous references in the section for numerical experiments. In order to simplify the implementation, we add the elliptic projector $\Pi_{h}^{\nabla}$ for all $v,w$ in the jump and average terms. In Section 4, we construct an enriching operator and present the corresponding error estimates. Section 5 develops the a posteriori analysis for the IPVEM, providing both upper and lower bounds. We do not include the jumps of the function itself in the computable posteriori error estimators despite the discontinuity of the interior penalty virtual element. Meanwhile, we make a comparison with the analysis of Morley-type virtual elements in [13]. An adaptive IPVEM is presented in Section 6 with several benchmark tests performed to illustrate the robustness of the residual-type a posteriori error estimator and demonstrate the efficiency of the adaptive IPVEM.
@@ 2 The continuous variational problem
$\left\{\begin{array}[]{ll}\Delta^{2}u=f&{\rm in}~{}~{}\Omega,\\ u=0,\quad\dfrac{\partial u}{\partial\bm{n}}=0&{\rm on}~{}~{}\partial\Omega,\end{array}\right.$ (2.1)
where $\bm{n}=(n_{1},n_{2})^{T}$ is the unit outer normal to $\partial\Omega$, $\Delta$ is the standard Laplacian operator. Throughout the article we assume $u\in H^{k+1}(\Omega)$ with $k\geq 2$.
@@
$\sum_{e\in\mathcal{E}_{h}}\int_{e}[vw]\mathrm{d}s=\sum_{e\in\mathcal{E}_{h}^{0}}\int_{e}[v]\{w\}\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}}\int_{e}\{v\}[w]\mathrm{d}s.$ (2.2)
Moreover, for any two quantities $a$ and $b$, “$a\lesssim b$” indicates “$a\leq Cb$” with the constant $C$ independent of the mesh size $h_{K}$, and “$a\eqsim b$” abbreviates “$a\lesssim b\lesssim a$”.
@@
where $a(u,v)=(\nabla^{2}u,\nabla^{2}v)$. To keep the flexibility of meshes, we will work on meshes satisfying the following assumption (cf. [10, 14]):
@@
As shown in [14], this condition covers the usual geometric assumptions frequently used in the context of VEMs. Under this geometric assumption, we can establish some fundamental results in VEM analysis as used in [21], with the constants hidden in the symbol $\lesssim$ depending only on the shape regularity and quasi-uniformity of the virtual triangulation ${\mathcal{T}_{K}}$.
According to the standard Dupont-Scott theory (cf. [7]), for all $v\in H^{l}(K)$ ($0\leq l\leq k$) there exists a certain $q\in\mathbb{P}_{l-1}(K)$ such that
@@
The following inequalities are very useful for our forthcoming analysis.
Lemma 2.1.
@@
  • 1.
    Trace inequality [7]:
    $\|v\|_{0,\partial K}\lesssim h_{K}^{1/2}|v|_{1,K}+h_{K}^{-1/2}\|v\|_{0,K},\quad v\in H^{1}(K),$ (2.5)
  • 2.
    Sobolev inequality [6]:
    $\|v\|_{\infty,K}\lesssim h_{K}|v|_{2,K}+|v|_{1,K}+h_{K}^{-1}\|v\|_{0,K},\quad v\in H^{2}(K),$ (2.6)
  • 3.
    Poincaré-Friedriches inequality [5]:
    $\|v\|_{0,K}\lesssim h_{K}|v|_{1,K}+\Big{|}\int_{\partial K}v\,{\rm d}s\Big{|},\quad v\in H^{1}(K).$ (2.7)
@@ 3 The interior penalty virtual element method
This section reviews the interior penalty virtual element space developed in [30].
@@ 3.1   The IP virtual element spaces
$\widetilde{V}_{k+2}(K)=\Big{\{}v\in H^{2}(K):\Delta^{2}v\in\mathbb{P}_{k}(K),v|_{e}\in\mathbb{P}_{k+2}(e),\partial_{\bm{n}}v|_{e}\in\mathbb{P}_{k+1}(e),e\subset\partial K\Big{\}},\quad k\geq 2.$
@@
Referring to [30, Lemma 3.4], we have the following lemma.
Lemma 3.1 (Inverse inequality).
For any $K\in\mathcal{T}_{h}$, for $m=1,2$ and $0\leq s\leq m$, there holds
$|v_{h}|_{m,K}\lesssim h_{K}^{s-m}|v_{h}|_{s,K},\quad v_{h}\in\widetilde{V}_{k+2}(K).$ (3.6)
@@
where $\mathcal{V}_{K}$ is the set of the vertices of $K$. By checking the right-hand side of the integration by parts formula
$(\nabla v_{h},\nabla q)_{K}=-(v_{h},\Delta q)_{K}+\sum\limits_{e\subset\partial K}\int_{e}v_{h}\partial_{\bm{n}_{e}}q\,\mathrm{d}s,$
one can find that this elliptic projection can be computed using the DoFs in (3.1)–(3.5). However, the goal of the IPVEM is to make $\Pi_{K}^{\nabla}v_{h}$ computable by only using the DoFs of $H^{1}$-conforming virtual element spaces given by (cf. [3, 1])
@@
To do so, Ref. [30] considered an approximation of the right-hand side by some numerical formula. The modified $H^{1}$-projection is defined as
@@
With the help of $\Pi_{K}^{\nabla}$, we can employ the standard enhancement technique by substituting the redundant DoFs of $v$ with those of $\Pi_{K}^{\nabla}v$ [1, 30]. The local interior penalty space is defined as
$V_{k}(K)=\{v\in\widetilde{V}_{k+2}(K):\tilde{\chi}^{g}(v)=\tilde{\chi}^{g}(\Pi_{K}^{\nabla}v),\tilde{\chi}^{n}(v)=\tilde{\chi}^{n}(\Pi_{K}^{\nabla}v),$
$\hskip 113.81102pt(v,q)_{K}=(\Pi_{K}^{\nabla}v,q)_{K}\quad q\in\mathbb{M}_{k}(K)\backslash\mathbb{M}_{k-2}(K)\}\quad k\geq 2,$
@@
Denote by $I_{h}:H^{2}(\Omega)\to V_{h}$ the interpolation operator. As shown in Lemma 3.11 of [30], we can derive the following interpolation error estimate.
Lemma 3.2.
For any $v\in H^{\ell}(K)$ with $12\leq\ell\leq k+$ 1, it holds
$|v-I_{h}v|_{m,K}\lesssim h_{K}^{\ell-m}|v|_{\ell,K},\quad K\in\mathcal{T}_{h},\quad m=0,1,2.$ (3.8)
@@
Lemma 3.3.
For every $v\in V_{k}(K)$, there hold
$|\Pi_{K}^{\nabla}v|_{m,K}\lesssim|v|_{m,K}\quad\mbox{and}\quad|\Pi_{K}^{\Delta}v|_{m,K}\lesssim|v|_{m,K},$
where $m=0,1,2$.
Proof.
The first inequality is presented in Corollary 3.7 of [30]. For the second one, it is simple to find that $|\Pi_{K}^{\Delta}v|_{2,K}\lesssim|v|_{2,K}$. For $m=1$, by the triangle inequality, the Poincaré-Friedrichs inequality (2.7), the boundedness of $|\Pi_{K}^{\Delta}v|_{2,K}$ and the inverse inequality (3.6), we derive
$|\Pi_{K}^{\Delta}v|_{1,K}$ $\leq|v-\Pi_{K}^{\Delta}v|_{1,K}+|v|_{1,K}\lesssim h_{K}|v-\Pi_{K}^{\Delta}v|_{2,K}+|v|_{1,K}$
$\lesssim h_{K}|v|_{2,K}+|v|_{1,K}\lesssim|v|_{1,K}.$
The case of $m=0$ can be deduced in the similar manner. ∎
3.2   The discrete IP bilinear form
Following the standard procedure in the VEMs (namely, we need to consider the computability, the $k$-consistency (3.17) and the stability (3.18)), we define the discrete bilinear form as
$a_{h}^{K}(v,w)=a^{K}(\Pi_{K}^{\Delta}v,\Pi_{K}^{\Delta}w)+h_{K}^{-2}S^{K}(v-\Pi_{K}^{\Delta}v,w-\Pi_{K}^{\Delta}w),\quad v,w\in V_{k}(K),$